# Training Set Reconstruction from Differentially Private Forests: How Effective is DP?

### v1

@@ Abstract
Recent research has shown that machine learning models are vulnerable to privacy attacks targeting their training data. Differential privacy (DP) has become a widely adopted countermeasure, as it offers rigorous privacy protections. In this paper, we introduce a reconstruction attack targeting state-of-the-art $\varepsilon$-DP random forests. By leveraging a constraint programming model that incorporates knowledge of the forest’s structure and DP mechanism characteristics, our approach formally reconstructs the most likely dataset that could have produced a given forest. Through extensive computational experiments, we examine the interplay between model utility, privacy guarantees, and reconstruction accuracy across various configurations. Our results reveal that random forests trained with meaningful DP guarantees can still leak substantial portions of their training data. Specifically, while DP reduces the success of reconstruction attacks, the only forests fully robust to our attack exhibit predictive performance no better than a constant classifier. Building on these insights, we provide practical recommendations for the construction of DP random forests that are more resilient to reconstruction attacks and maintain non-trivial predictive performance.
Machine Learning, ICML
@@ 1 Introduction
The success of machine learning algorithms in critical applications, such as medicine, kidney attribution and credit scoring (Dildar et al., 2021; Aziz et al., 2021; Dastile et al., 2020), hinges on access to large volumes of sensitive data, including medical records, personal histories and financial information. This extensive data is crucial for training models capable of achieving high accuracy. At the same time, trustworthiness in these applications demands transparency, as emphasized by regulations like Article 13 of the EU AI Act [^1] https://artificialintelligenceact.eu/article/13/. Transparency involves the ability to audit models and explain their outputs and inner workings, which fosters public confidence and ensures compliance with legal requirements.
@@
However, transparency can inadvertently create vulnerabilities. When models are made public, even via black-box APIs, they become susceptible to privacy attacks that can infer sensitive information about their training data (Rigaki & Garcia, 2023). Recent studies on reconstruction attacks (Dwork et al., 2017) demonstrated that large language models (Carlini et al., 2021), diffusion models (Carlini et al., 2023), neural networks (Haim et al., 2022), random forests (RFs) and a wide range of interpretable models (Ferry et al., 2024a, b) inherently leak part of their training data through their structure or outputs.
@@
To mitigate such privacy risks, differential privacy (DP) (Dwork & Roth, 2014) has emerged as the de facto standard, as it limits the amount of information an adversary can gain regarding any single individual in the dataset. DP theoretically bounds the success of membership inference attacks (which aim at inferring the absence or presence of a given individual in a model’s training data), but also empirically provides a form of protection against other attacks. For instance, recent work on deep learning for medical imaging (Ziller et al., 2024) has shown that DP mitigates the success of some reconstruction attacks, even for loose privacy guarantees. However, DP imposes significant trade-offs as ensuring privacy involves adding noise or other modifications that degrade a model’s predictive performance. Consequently, practitioners must carefully balance privacy guarantees and model utility in sensitive domains.
@@
In this paper, we focus on reconstruction attacks targeting tree ensembles, such as RFs. These models frequently achieve state-of-the-art performance on tabular data, but recent research (Ferry et al., 2024b) has revealed that their structure can inherently expose (nearly) all of their training data. While several DP tree-based models have been designed (Fletcher & Islam, 2019), it remains unclear whether DP can effectively prevent such privacy leaks. To address this gap, the main contributions of our study are as follows:
@@
  • We introduce an effective reconstruction attack against state-of-the-art DP RFs. Building on the attack paradigm pioneered by (Ferry et al., 2024b), our approach frames the reconstruction attack as the solution of a Constraint Programming (CP) model, which identifies a most likely dataset capable of generating the observed forest and noisy counts. We demonstrate that this method can retrieve significant portions of RFs’ training data, even under tight privacy budgets.
@@
  • Through extensive experiments, we show that DP RFs, while formally satisfying privacy guarantees, can still leak substantial portions of their training data. In particular, our findings reveal that all the considered DP RFs with non-trivial predictive performance remain vulnerable to our reconstruction attack. The source code of our method, permitting the reproducibility of all the experiments, is openly available at https://github.com/vidalt/DRAFT-DP, under a MIT license.
@@
  • We provide key insights and practical recommendations regarding the factors influencing the trade-offs between DP RFs’ predictive performance and their vulnerability to reconstruction attacks. These insights aim to guide practitioners in effectively deploying DP RFs in real-world scenarios.
@@ 2.1 Differential Privacy
Originally formalized in Dwork (2006), DP enables the computation of useful statistics on private data while strictly limiting the information that can be inferred about any single individual. A randomized mechanism $\mathcal{K}$ satisfies $\varepsilon$-DP if, for any pair of datasets $D_{1}$ and $D_{2}$ differing by at most one record, and all $U\subseteq\text{Range}(\mathcal{K})$:
@@
Let $q$ represent a deterministic query such as counts, sums, averages, or, in the context of tree ensembles, the number of samples of a particular class in a data subset. DP can give protected access to this query through a probabilistic mechanism $\mathcal{K}$, which outputs $q$ with additional noise. As detailed in Dwork et al. (2016), the amount of noise needed to protect the data depends on the global sensitivity of the query, defined as $S(q)=\max||q(D_{1})-q(D_{2})||_{1},$ in which the maximum is taken over all pairs of datasets $D_{1}$ and $D_{2}$ differing in at most one element. Thus, $S(q)$ quantifies the largest possible change in the query output due to a single record’s addition or removal (e.g., $S(q)=1$ for counts). For instance, the Laplace mechanism, a widely used DP method, adds noise sampled from a Laplace distribution:
@@
When multiple DP mechanisms are combined, the overall privacy budget is governed by two fundamental composition theorems (McSherry & Talwar, 2007). First, sequential composition states that applying several $\varepsilon_{i}$-DP mechanisms to the same dataset results in $\sum_{i}\varepsilon_{i}$-DP. In contrast, parallel composition states that applying several $\varepsilon_{i}$-DP mechanisms to disjoint subsets of data results in $\max_{i}\varepsilon_{i}$-DP. Finally, the postprocessing property ensures that any function applied to the output of $\mathcal{K}$ does not affect the DP guarantee, providing flexibility for downstream analysis.
DP mechanisms have been widely adopted to protect privacy when releasing machine learning models. For example, noise can be added to gradients during neural network training (Abadi et al., 2016), or split thresholds can be perturbed in decision tree predictors (Fletcher & Islam, 2019). In particular, DP can mitigate privacy risks such as membership inference or model inversion (Ye et al., 2022), and implementations of DP mechanisms have been integrated within libraries, such as IBM’s diffprivlib (Holohan et al., 2019) or Google’s differential-privacy [^2] https://github.com/google/differential-privacy/.
@@ 2.2 Differentially-Private Random Forests
We consider a training set $D={\{\mathbf{x}_{k};c_{k}\}}^{N}_{k=1}$ in which each example $k$ is characterized by a vector $\mathbf{x}_{k}\in\{0,1\}^{M}$ of $M$ binary attributes and a class $c_{k}\in\mathcal{C}$. Categorical features can be transformed into this format using one-hot encoding, as is standard practice in tree ensemble training, while continuous features can be discretized through binning. Based on this dataset, a random forest $\mathcal{T}$ is built as an ensemble of decision trees $t\in\mathcal{T}$. Each tree consists of a set of internal nodes $\smash{\mathcal{V}_{t}^{I}}$, in which each node $v\in\smash{\mathcal{V}_{t}^{I}}$ contains a binary condition on the value of an attribute, and a set of leaves $\smash{\mathcal{V}_{t}^{L}}$. When an example is evaluated, it traverses the tree by descending to the left child $l(v)$ if the condition is satisfied, or to the right child $r(v)$ otherwise. Once a leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ (terminal node) is reached, it contributes to the class distribution associated with that leaf, which is determined by the per-class counts of training examples reaching the leaf.
In most random forest implementations, including popular libraries such as scikit-learn (Pedregosa et al., 2011), predictions are made using soft voting: each tree outputs class probabilities based on the normalized class counts in its leaves and the ensemble prediction averages these probabilities. Specifically, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$, the model stores $n_{tvc}$, the number of training examples of class $c$ that reach $v$. This information is essential for on-device inference and audits but also introduces risks, as it exposes model parameters to potential adversaries (Ferry et al., 2024b).
To address these risks, an extensive body of research has focused on constructing DP RFs to conciliate model accessibility and privacy protection with many of these seminal methods being surveyed by Fletcher & Islam (2019). Among these, the method proposed by Jagannathan et al. (2012), which combines randomized construction and noisy counts publication, was the first to achieve strong empirical predictive performance and formal DP guarantees. Consequently, we use it a foundation of our experimental study.
In this approach, each tree is trained on the complete dataset $D$. Each tree of a chosen depth $d$ is grown by randomly selecting attributes and split values for all internal nodes $v\in\smash{\mathcal{V}_{t}^{I}}$, avoiding direct access to $D$ during this phase to preserve the DP budget. Subsequently, the class counts $n_{tvc}$ at each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are obfuscated using the Laplace mechanism (Dwork et al., 2016) and cast to integer values, as illustrated in Figure 1.
@@
Let $\varepsilon$ denote the total privacy budget. Since all trees $t\in\mathcal{T}$ are trained on the same dataset $D$, sequential composition applies, dividing the budget across trees such that each tree has a budget of $\varepsilon_{t}=\varepsilon/|\mathcal{T}|$. Since the attributes and split values within each tree are chosen randomly without accessing the training data, the noisy computation of class counts is the sole operation consuming the DP budget. Within each tree, parallel composition applies since the leaves have disjoint supports. Similarly, within a leaf, the counting operations for each class are independent, also satisfying parallel composition. Accordingly, the DP budget allocated to publish each class count $n_{tvc}$ for $t\in\mathcal{T},v\in\smash{\mathcal{V}_{t}^{L}},c\in\mathcal{C}$ is $\varepsilon_{v}=\varepsilon_{t}=\varepsilon/|\mathcal{T}\rvert$. Since class counts have a global sensitivity $S(q)=1$, the noisy class counts are computed as:
@@ 3 Reconstruction Attack against DP RFs
We assume white-box access to an $\varepsilon$-DP RF constructed using the algorithm of Jagannathan et al. (2012), as described in Section 2.2. This includes access to the structure of the forest (split attributes and values) and the (noisy) counts at the leaves. Consistent with the literature on reconstruction attacks (e.g., Dwork et al., 2017; Ferry et al., 2024b), we also assume knowledge of the number of attributes $M$ and training examples $N$. While the knowledge of $M$ is necessary for inference, the knowledge of $N$ simplifies the attack but is not strictly required. In particular, in Appendix A, we demonstrate that our attack can reconstruct the dataset without prior knowledge of $N$, making it the first method to target full dataset reconstruction without this assumption. We also assume the privacy budget $\varepsilon$ is known. This does not degrade the DP guarantees, is generally disclosed by practitioners [^3] For example, Desfontaines (2021)’s website documents real-world applications of DP and their associated budgets., and researchers have advocated for its transparency (Dwork et al., 2019).
Reconstruction Approach.
@@
Our approach is grounded on the CP formulation introduced in Ferry et al. (2024b), with additional constraints, variables and an objective specifically designed to model the (observed) noisy counts, the underlying (guessed) true counts and their likelihood. Solving the original problem formulation (called DRAFT) was shown to be NP-Hard, though practical at scale with state-of-the-art CP solvers (Perron & Didier, 2024). As our reconstruction problem for DP RFs generalizes it (i.e., DRAFT is a special case with no noise), solving it is also NP-Hard.
@@
Our formulation revolves around three main groups of decision variables. The objective of a CP solver is to find compatible variable values that satisfy a given set of constraints, yielding a feasible solution. Moreover, as explained later in this section, an objective function steering the search towards likely solutions (reconstructions) will be included. First, for each tree $t\in\mathcal{T}$, the (guessed) true number of examples of class $c\in\mathcal{\mathcal{C}}$ in each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are encoded using a variable $n_{tvc}\in\mathbb{N}$. Next, a set of variables encodes each reconstructed example $k\in\{1,\dots,N\}$: $x_{ki}\in\{0,1\}$ encodes the value of its $i^{\text{th}}$ attribute $i\in\{1,\dots,M\}$, while $z_{kc}\in\{0,1\}$ indicates if it belongs to class $c\in\mathcal{C}$. Third, the path of each example $k\in\{1,\dots,N\}$ through each tree $t\in\mathcal{T}$ is modelled through a binary variable $y_{tvkc}\in\{0,1\}$, which is $1$ if it is classified by leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ as class $c\in\mathcal{C}$, and 00 otherwise. The following constraints linking these variables are then enforced:
$\sum\limits_{v\in\smash{\mathcal{V}_{t}^{L}}}\sum\limits_{c\in\mathcal{C}}n_{tvc}=N$ $t\in\mathcal{T}$ (1)
$\sum\limits_{c\in\mathcal{C}}z_{kc}=1$ $k\in\{1,\dots,N\}$ (2)
$z_{kc}=\sum\limits_{\smash{\mathclap{v\in\smash{\mathcal{V}_{t}^{L}}}}}y_{tvkc}$ $k\in\{1,\dots,N\},t\in\mathcal{T},c\in\mathcal{C}$ (3)
$\sum_{k=1}^{\smash{N}}y_{tvkc}=n_{tvc}$ $t\in\mathcal{T},~{}v\in\smash{\mathcal{V}_{t}^{L}},~{}c\in\mathcal{C}$ (4)
@@
In a nutshell, Constraint 1 ensures that the example counts sum up to $N$ within each tree. Constraint 2 guarantees that each example is associated to exactly one class. Constraint 3 enforces that each example contributes only to the counts of its assigned class. Moreover, Constraint 4 ensures consistency between the counts at each leaf and the examples assigned to it. Finally, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ of each tree $t\in\mathcal{T}$, let $\Phi_{v}^{+}$ (respectively $\Phi_{v}^{-}$) be the set of binary attributes that must equal $1$ (respectively 00) for an example to fall into $v$. Then, the following constraint enforces consistency between the examples’ assignments to the leaves, and the value of their attributes:
@@
$\sum\limits_{c\in\mathcal{C}}y_{tvkc}=1\Rightarrow$ $\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{+}}x_{ki}=1\right)\wedge\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{-}}x_{ki}=0\right)$
$\quad t\in\mathcal{T},~{}v\in\smash{\mathcal{V}_{t}^{L}},~{}k\in\{1,\dots,N\}.$ (5)
@@
The aforementioned variables and constraints define the domain of all possible datasets and associated counts $n_{tvc}$ compatible with the structure and split values of an observed forest. Among all solutions in this domain, we aim at retrieving the one that is the most likely to have generated the observed noisy counts $n_{tvc}^{*}$ through the DP mechanism. To this end, we formulate through Constraint 6 the noise $\Delta_{tvc}$ that must have been added to each count to achieve the observed values.
@@
We know that these noise values correspond to i.i.d realizations of random variables $Y_{tvc}$ from a Laplace distribution $\text{Lap}(1/\varepsilon_{v})$ cast to integer. Given this information, the likelihood of a complete solution associated to a set of $\Delta_{tvc}$ values is:
$\mathcal{P}_{\boldsymbol{\Delta}}=\prod\limits_{t,v,c}\mathbb{P}(\textrm{int}(Y_{tvc})=\Delta_{tvc}).$ (7)
Maximizing this likelihood is equivalent to maximizing its logarithm, leading to the following objective function:
$\log(\mathcal{P}_{\boldsymbol{\Delta}})$ $=\max\sum\limits_{t,v,c}\log(\mathbb{P}(\textrm{int}(Y_{tvc})=\Delta_{tvc})).$ (8)
@@
In practice, we pre-compute constants $p_{l}$ for all values $l\in\{-\delta,\dots,\delta\}$ such that $\mathbb{P}(\textrm{int}(Y_{tvc})\in\{-\delta,\dots,\delta\})\geq 0.999$. Details on the computation of $\delta$ are provided in Appendix B, but typically, $\delta=\lceil 12/\varepsilon_{v}\rceil$ is sufficient.
With this, we can additionally restrict the feasible domain of the variables to $n_{tvc}\in\{\max(0,n_{tvc}^{*}-\delta),\dots,n_{tvc}^{*}+\delta\}$ and $\Delta_{tvc}\in\{-\delta,\dots,\delta\}$ and formulate the objective as:

### v2

@@ Abstract
Recent research has shown that machine learning models are vulnerable to privacy attacks targeting their training data. To mitigate these risks, differential privacy (DP) has become a widely adopted countermeasure, as it offers rigorous privacy protection. In this paper, we introduce a reconstruction attack targeting state-of-the-art $\varepsilon$-DP random forests. By leveraging a constraint programming model that incorporates knowledge of the forest’s structure and DP mechanism characteristics, our approach formally reconstructs the most likely dataset that could have produced a given forest. Through extensive computational experiments, we examine the interplay between model utility, privacy guarantees and reconstruction accuracy across various configurations. Our results reveal that random forests trained with meaningful DP guarantees can still leak portions of their training data. Specifically, while DP reduces the success of reconstruction attacks, the only forests fully robust to our attack exhibit predictive performance no better than a constant classifier. Building on these insights, we provide practical recommendations for the construction of DP random forests that are more resilient to reconstruction attacks and maintain non-trivial predictive performance.
@@ 1 Introduction
The success of machine learning algorithms in critical applications, such as medicine, kidney attribution and credit scoring [Dildar et al., 2021, Aziz et al., 2021, Dastile et al., 2020], hinges on access to large volumes of sensitive data, including medical records, personal histories and financial information. This extensive data is crucial for training models capable of achieving high accuracy. At the same time, trustworthiness in these applications demands transparency, as emphasized by regulations like Article 13 of the EU AI Act [^1] https://artificialintelligenceact.eu/article/13/. Transparency involves the ability to audit models and explain their outputs and inner workings, which fosters public confidence and ensures compliance with legal requirements.
@@
However, transparency can inadvertently create vulnerabilities. When models are made public, even via black-box APIs, they become susceptible to privacy attacks that can infer sensitive information about their training data [Rigaki and Garcia, 2023]. Recent studies on reconstruction attacks [Dwork et al., 2017] demonstrated that large language models [Carlini et al., 2021], diffusion models [Carlini et al., 2023], neural networks [Haim et al., 2022], random forests (RFs) and a wide range of interpretable models [Ferry et al., 2024a, b] inherently leak part of their training data through their structure or outputs.
@@
To mitigate such privacy risks, differential privacy (DP) [Dwork and Roth, 2014] has emerged as the de facto standard, as it limits the amount of information an adversary can gain regarding any single individual in the dataset. DP theoretically bounds the success of membership inference attacks (which aim at inferring the absence or presence of a given individual in a model’s training data), but also empirically provides a form of protection against other attacks. For instance, recent work on deep learning for medical imaging [Ziller et al., 2024] has shown that DP mitigates the success of some reconstruction attacks, even for loose privacy guarantees. However, DP imposes significant trade-offs as ensuring privacy involves adding noise or other modifications that degrade a model’s predictive performance. Consequently, practitioners must carefully balance privacy guarantees and model utility in sensitive domains.
@@
In this paper, we focus on reconstruction attacks targeting tree ensembles, such as RFs. These models frequently achieve state-of-the-art performance on tabular data, but recent research [Ferry et al., 2024b] has revealed that their structure can inherently expose their training data. While several DP tree-based models have been designed [Fletcher and Islam, 2019], it remains unclear how well DP can effectively prevent such privacy leaks. To address this gap, the main contributions of our study are:
@@
  • We introduce an effective reconstruction attack against state-of-the-art DP RFs. Building on the attack paradigm pioneered by [Ferry et al., 2024b], our approach frames the reconstruction attack as a Constraint Programming (CP) model, which identifies a most likely dataset capable of generating the observed RF and noisy counts.
@@
  • Through extensive experiments, we show that DP RFs, while formally satisfying privacy guarantees, can still be exploited to retrieve portions of their training data. In particular, all investigated DP RFs with non-trivial predictive performance reveal information about their training data, which often extends beyond distributional patterns to specificities of the dataset used. To enable the reproducibility of all the experiments, the source code of our method is openly available at https://github.com/vidalt/DRAFT-DP, under a MIT license.
@@
  • We provide key insights regarding the factors influencing the trade-offs between DP RFs’ predictive performance and their vulnerability to reconstruction attacks. These insights aim to guide practitioners in effectively deploying DP RFs in real-world scenarios.
@@ 2.1 Differential Privacy
Originally formalized in Dwork [2006], DP enables the computation of useful statistics on private data while strictly limiting the information that can be inferred about any single individual. A randomized mechanism $\mathcal{K}$ satisfies $\varepsilon$-DP if, for any pair of datasets $D_{1}$ and $D_{2}$ differing by at most one record, and all $U\subseteq\text{Range}(\mathcal{K})$:
@@
Let $q$ represent a deterministic query such as counts, sums, averages, or, in the context of tree ensembles, the number of samples of a particular class in a data subset. DP can give protected access to this query through a probabilistic mechanism $\mathcal{K}$, which outputs $q$ with additional noise. As detailed in Dwork et al. [2016], the amount of noise needed to protect the data depends on the global sensitivity of the query, defined as $S(q)=\max||q(D_{1})-q(D_{2})||_{1},$ in which the maximum is taken over all pairs of datasets $D_{1}$ and $D_{2}$ differing in at most one element. Thus, $S(q)$ quantifies the largest possible change in the query output due to a single record’s addition or removal (e.g., $S(q)=1$ for counts). For instance, the Laplace mechanism, a widely used DP method, adds noise sampled from a Laplace distribution:
@@
When multiple DP mechanisms are combined, the overall privacy budget is governed by two fundamental composition theorems [McSherry and Talwar, 2007]. First, sequential composition states that applying several $\varepsilon_{i}$-DP mechanisms to the same dataset results in $\sum_{i}\varepsilon_{i}$-DP. In contrast, parallel composition states that applying several $\varepsilon_{i}$-DP mechanisms to disjoint subsets of data results in $\max_{i}\varepsilon_{i}$-DP. Finally, the postprocessing property ensures that any function applied to the output of $\mathcal{K}$ does not affect the DP guarantee, providing flexibility for downstream analysis.
DP mechanisms have been widely adopted to protect privacy when releasing machine learning models. For example, noise can be added to gradients during neural network training [Abadi et al., 2016], or split thresholds can be perturbed in decision tree predictors [Fletcher and Islam, 2019]. In particular, DP can mitigate privacy risks such as membership inference or model inversion [Ye et al., 2022], and implementations of DP mechanisms have been integrated within popular libraries, such as IBM’s diffprivlib [Holohan et al., 2019] or Google’s differential-privacy [Google, ].
@@ 2.2 Differentially-Private Random Forests
We consider a training set $D={\{\mathbf{x}_{k};c_{k}\}}^{N}_{k=1}$ in which each example $k$ is characterized by a vector $\smash{\mathbf{x}_{k}\in\{0,1\}^{M}}$ of $M$ binary attributes and a class $c_{k}\in\mathcal{C}$. Categorical features can be transformed into this format using one-hot encoding, as is standard practice in tree ensemble training, while continuous features can be discretized through binning. Based on this dataset, a random forest $\mathcal{T}$ is built as an ensemble of decision trees $t\in\mathcal{T}$. Each tree consists of a set of internal nodes $\smash{\mathcal{V}_{t}^{I}}$, in which each node $v\in\smash{\mathcal{V}_{t}^{I}}$ contains a binary condition on the value of an attribute, and a set of leaves $\smash{\mathcal{V}_{t}^{L}}$. When an example is evaluated, it traverses the tree by descending to the left child $l(v)$ if the condition is satisfied, or to the right child $r(v)$ otherwise. Once a leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ (terminal node) is reached, it contributes to the class distribution associated with that leaf, which is determined by the per-class counts of training examples reaching the leaf.
In most random forest implementations, including popular libraries such as scikit-learn [Pedregosa et al., 2011], predictions are made using soft voting: each tree outputs class probabilities based on the normalized class counts in its leaves and the ensemble prediction averages these probabilities. Specifically, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ in tree $t$, the model stores $n_{tvc}$, the number of training examples of class $c$ that reach $v$. This information is essential for on-device inference and audits but also introduces risks, as it exposes model parameters to potential adversaries [Ferry et al., 2024b].
To address these risks, an extensive body of research has focused on constructing DP RFs to conciliate model accessibility and privacy protection, with many of these seminal methods being surveyed by Fletcher and Islam [2019]. Among these, the method proposed by Jagannathan et al. [2012], which combines randomized tree structure and noisy counts publication, was the first to achieve strong empirical predictive performance and formal DP guarantees. Consequently, we use it as a foundation of our experimental study. In this approach, each tree is trained on the complete dataset $D$. Each tree of a chosen depth $d$ is grown by randomly selecting attributes and split values for all internal nodes $v\in\smash{\mathcal{V}_{t}^{I}}$, avoiding direct access to $D$ during this phase to preserve the DP budget. Subsequently, the class counts $n_{tvc}$ at each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are obfuscated using the Laplace mechanism [Dwork et al., 2016] and cast to integer values, as illustrated in Figure 1.
@@
Let $\varepsilon$ denote the total privacy budget. Since all trees $t\in\mathcal{T}$ are trained on the same dataset $D$, sequential composition applies, such that each tree has a budget of $\varepsilon_{t}=\varepsilon/|\mathcal{T}|$. Since the attributes and split values within each tree are chosen randomly without accessing the training data, the noisy computation of class counts is the sole operation consuming the DP budget. Within each tree, parallel composition applies since the leaves have disjoint supports. Similarly, within a leaf, the counting operations for each class are independent, also satisfying parallel composition.
Accordingly, the DP budget allocated to publish each class count $n_{tvc}$ for $t\in\mathcal{T},v\in\smash{\mathcal{V}_{t}^{L}},c\in\mathcal{C}$ is $\varepsilon_{v}=\varepsilon_{t}=\varepsilon/|\mathcal{T}\rvert$. Since class counts have a global sensitivity $S(q)=1$, the noisy class counts are computed as:
@@ 3 Reconstruction Attack against DP RFs
Attack Model. We assume white-box access to an $\varepsilon$-DP RF constructed using the algorithm of Jagannathan et al. [2012], as described in Section 2.2. This includes access to the structure of the forest (split attributes and values) and the (noisy) counts at the leaves. Consistent with the literature on reconstruction attacks [e.g., Dwork et al., 2017, Ferry et al., 2024b], we also assume knowledge of the number of attributes $M$ and training examples $N$. While the knowledge of $M$ is necessary for inference, the knowledge of $N$ simplifies the attack but is not strictly required. In particular, in Appendix A, we demonstrate that our attack can reconstruct the dataset without prior knowledge of $N$, making it the first method to target full dataset reconstruction without this assumption. We also assume the privacy budget $\varepsilon$ is known. This does not degrade the DP guarantees, is generally disclosed by practitioners [^2] For example, Desfontaines [2021] documents real-world applications of DP and their associated budgets., and researchers have advocated this for transparency [Dwork et al., 2019].
@@
Reconstruction Approach. Our approach is grounded on the CP formulation introduced in Ferry et al. [2024b], with additional constraints, variables, and an objective specifically designed to model the (observed) noisy counts, the underlying (guessed) true counts, and their likelihood. Solving the original problem formulation (called DRAFT) was shown to be NP-Hard. Our reconstruction problem for DP RFs can be seen as a generalization (i.e., DRAFT is a special case with no noise) and is also NP-hard, though scalable with state-of-the-art CP solvers [Perron and Didier, 2024].
@@
Our formulation revolves around three main groups of decision variables. The objective of a CP solver is to find compatible variable values that satisfy a given set of constraints, yielding a feasible solution. Moreover, as explained later in this section, an objective function steering the search towards likely solutions (reconstructions) will be included. First, for each tree $t\in\mathcal{T}$, the (guessed) true number of examples of class $c\in\mathcal{\mathcal{C}}$ in each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ are encoded using a variable $n_{tvc}\in\mathbb{N}$. Next, a set of variables encodes each reconstructed example $k\in\{1,\dots,N\}$: $x_{ki}\in\{0,1\}$ encodes the value of its $i^{\text{th}}$ attribute $i\in\{1,\dots,M\}$, while $z_{kc}\in\{0,1\}$ indicates if it belongs to class $c\in\mathcal{C}$. Third, the path of each example $k\in\{1,\dots,N\}$ through each tree $t\in\mathcal{T}$ is modelled using a binary variable $y_{tvkc}\in\{0,1\}$, which is $1$ if it is classified by leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ as class $c\in\mathcal{C}$, and 00 otherwise. The following constraints linking these variables are then enforced: $\sum\limits_{v\in\smash{\mathcal{V}_{t}^{L}}}\sum\limits_{c\in\mathcal{C}}n_{tvc}=N$ $t\in\mathcal{T}$ (1) $\sum\limits_{c\in\mathcal{C}}z_{kc}=1$ $k\in\{1,\dots,N\}$ (2) $z_{kc}=\sum\limits_{\smash{\mathclap{v\in\smash{\mathcal{V}_{t}^{L}}}}}y_{tvkc}$ $k\in\{1,\dots,N\},t\in\mathcal{T},c\in\mathcal{C}$ (3) $\sum_{k=1}^{\smash{N}}y_{tvkc}=n_{tvc}$ $t\in\mathcal{T},~{}v\in\smash{\mathcal{V}_{t}^{L}},~{}c\in\mathcal{C}$ (4)
@@
In a nutshell, Constraint (1) ensures that the example counts sum up to $N$ within each tree. Constraint (2) guarantees that each example is associated to exactly one class. Constraint (3) enforces that each example contributes only to the counts of its assigned class. Moreover, Constraint (4) ensures consistency between the counts at each leaf and the examples assigned to it. Finally, for each leaf $v\in\smash{\mathcal{V}_{t}^{L}}$ of each tree $t\in\mathcal{T}$, let $\Phi_{v}^{+}$ (respectively $\Phi_{v}^{-}$) be the set of binary attributes that must equal $1$ (respectively 00) for an example to fall into $v$. Then, the following constraint enforces consistency between the examples’ assignments to the leaves, and the value of their attributes:
@@
$\sum\limits_{c\in\mathcal{C}}y_{tvkc}=1\Rightarrow$ $\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{+}}x_{ki}=1\right)\wedge\left(\displaystyle\bigwedge_{i\in\Phi_{v}^{-}}x_{ki}=0\right)$ $t\in\mathcal{T},~{}v\in\smash{\mathcal{V}_{t}^{L}},~{}k\in\{1,\dots,N\}$ (5)
@@
The aforementioned variables and constraints define the domain of all possible datasets and associated counts $n_{tvc}$ compatible with the structure and split values of an observed forest. Among all solutions in this domain, we aim at retrieving the one that is the most likely to have generated the observed noisy counts $n_{tvc}^{*}$ through the DP mechanism. To this end, we formulate through Constraint (6) the noise $\Delta_{tvc}$ that must have been added to each count to achieve the observed values.
@@
We know that these noise values correspond to i.i.d realizations of random variables $Y_{tvc}$ from a Laplace distribution $\text{Lap}(1/\varepsilon_{v})$ cast to integer. Given this information, the likelihood of a complete solution associated to a set of $\Delta_{tvc}$ values is defined in Equation (7). Maximizing this likelihood is equivalent to maximizing its logarithm, leading to the objective function in Equation (8). $\mathcal{P}_{\boldsymbol{\Delta}}=\prod\limits_{t,v,c}\mathbb{P}(\textrm{int}(Y_{tvc})=\Delta_{tvc})$ (7) $\max\log(\mathcal{P}_{\boldsymbol{\Delta}})=\sum\limits_{t,v,c}\log(\mathbb{P}(\textrm{int}(Y_{tvc})=\Delta_{tvc}))$ (8)
@@
In practice, we pre-compute constants $p_{l}$ for all values $l\in\{-\delta,\dots,\delta\}$ such that $\mathbb{P}(\textrm{int}(Y_{tvc})\in\{-\delta,\dots,\delta\})\geq 0.999$. Details on the computation of $\delta$ are provided in Appendix B, but typically, $\delta=\lceil 12/\varepsilon_{v}\rceil$ is sufficient. With this, we can additionally restrict the feasible domain of the variables to $n_{tvc}\in\{\max(0,n_{tvc}^{*}-\delta),\dots,n_{tvc}^{*}+\delta\}$ and $\Delta_{tvc}\in\{-\delta,\dots,\delta\}$ and formulate the objective as: